OpenStudy (anonymous):

The height of an equilateral triangular prism increases by three units. The new volume is more than the original by how much? Answer three times the length of the height three times the area of the base three more than the the area of the base three more than the length of the height

OpenStudy (phi):

Can you answer, what is the formula for the volume of a prism with a height of 3?

Directrix (directrix):

three times the area of the base ---------------------------- Original Prism Volume = Bh where B is the area of the base and h is the height of the prism. New Prism Volume = B(h+3) where the height of the original prism has been increased by 3.. V = B(h+3) V = Bh + 3B ---> compare to V = Bh